Overhead Variances, Two- And Three-Variance Analyses

Oerstman, Inc., uses a standard costing system and develops its overhead rates from the current annual budget. The budget is based on an expected annual output of 122,000 units requiring 488,000 direct labor hours. (Practical capacity is 508,000 hours.) Annual budgeted overhead costs total $746,640, of which $541,680 is fixed overhead. A total of 119,300 units using 486,000 direct labor hours were produced during the year. Actual variable overhead costs for the year were $241,600, and actual fixed overhead costs were $555,450.

Required:

1. Compute overhead variances using a two-variance analysis.

Budget Variance $
Volume Variance $

2. Compute overhead variances using a three-variance analysis.

Spending Variance $
Efficiency Variance $
Volume Variance $
Answer #1

finding of variance using output method;

budgeted output=122000 units

actual output=119300 units

total budgeted overheads =746640

actual overheads= fixed+variable=555450+241600=797050

standard rate = budgeted cost/budgeted output=746640/122000=6.12

actual rate = actual cost/actual output=797050/119200=6.687

as per two variance analysis

budgeted variance = standard rate*budgeted units-actual rate*budgeted units

                            =746640-797050=50410(adverse)

volume variance=standard rate(actual-budgeted output)

                       = 6.12(119200-122000)=17136(adverse)

as per three variance;

budgeted variance = standard rate*budgeted units-actual rate*budgeted units

                            =746640-797050=50410(adverse)

volume variance=17136-13770=3366(adverse)

efficiency variance=13770(adverse) i.e... actual fixed cost-budgeted fixed cost